# Multimodal Foundation Models for Early Disease Detection

### v1

@@ Abstract
Healthcare generates diverse streams of data, including electronic health records (EHR), medical imaging, genetics, and ongoing monitoring from wearable devices. Traditional diagnostic models frequently analyze these sources in isolation, which constrains their capacity to identify cross-modal correlations essential for early disease diagnosis. Our research presents a multimodal foundation model that consolidates diverse patient data through an attention-based transformer framework. At first, dedicated encoders put each modality into a shared latent space. Then, they combine them using multi-head attention and residual normalization. The architecture is made for pretraining on many tasks, which makes it easy to adapt to new diseases and datasets with little extra work. We provide an experimental strategy that uses benchmark datasets in oncology, cardiology, and neurology, with the goal of testing early detection tasks. The framework includes data governance and model management tools in addition to technological performance to improve transparency, reliability, and clinical interpretability. The suggested method works toward a single foundation model for precision diagnostics, which could improve the accuracy of predictions and help doctors make decisions.
@@ I Introduction
The swift aggregation of manifold patient data poses both considerable analytical hurdles and unique opportunities. Healthcare data today come in many modalities, ranging from imaging such as MRI and CT scan, to sequential streams like wearable sensor outputs and electronic health records (EHRs). They also include audio (heart sounds, breathing patterns, or interview recordings), written sources (clinical documentation and scientific articles), video (for example, surgical recordings), as well as molecular-level information such as genomics and proteomics [1]. EHR further capture a timeline of clinical events, where variables such as diagnoses, medications, and procedures are captured at each visit . When combined, these sequences provide a detailed record of an individual patient’s medical history across multiple encounters [2] [3].
As healthcare providers and researchers aim to improve patient care through precision medicine [4], clinical decision-making significantly depends on the synthesis of these information sources. However, current predictive models used for this purpose are typically limited to a single data modality, which constrains their ability to provide comprehensive, patient-centric predictions as they could not catch the intricate interdependencies that exist across these heterogeneous data types. This limitation reduces the potential for genuinely patient-centric insights and constrains predictive accuracy.
In recent years, as deep learning has been producing models and tools capable of capturing complex relationships across diverse data modalities, attention-based architectures in particular, have shown the ability to dynamically evaluate the relevance of different inputs, allowing models to emphasize the information most critical for a given patient or predictive task. When pretrained on large, heterogeneous datasets, such models learn generalizable representations that can be adapted to multiple downstream diagnostic applications [5]. This paradigm, well established in natural language processing and computer vision [6], positions our approach as a foundation model for healthcare, providing reusable multimodal representations to support early disease detection and other clinical tasks. Despite these advancements, there are still many challenges because multimodal datasets are often fragmented, and individual patients may have partial or missing modalities [7]. Additionally, high-dimensional data, like whole-genome sequences or high-resolution images, make model integration and training even more difficult. Furthermore, models must provide transparent reasoning to support physician trust and facilitate actionable decisions; interpretability is a crucial requirement for clinical adoption.
The attention-based transformer architecture has proven highly effective for large-scale pretraining, where it learns rich contextual representations that enhance performance on a variety of downstream tasks, including sentiment analysis, information retrieval, and entity recognition [8]. Building on this paradigm, foundation models trained on massive and diverse datasets extend these benefits by enabling stronger contextual reasoning, broader generalization, and emergent prompting capabilities during inference [9]. That is why we propose a transformer-based foundation model architecture to resolve user challenges, which is intended to integrate EHR, imaging, genomic, and wearable sensor data. Our method utilizes cross-modal attention mechanisms to identify interdependencies between modalities and to facilitate pretraining on a variety of tasks, thereby facilitating the development of generalizable, patient-specific representations. This positions our approach as a multimodal foundation model, capable of serving as a base for diverse clinical prediction and diagnostic tasks. This work endeavors to facilitate the advancement of integrated diagnostic tools that are capable of enhancing patient outcomes and supporting precision medicine initiatives by integrating heterogeneous clinical data streams within a unified computational framework.
@@ II-A Multimodal Integration
The wide use of EHR systems has led to the development of predictive models that can support clinical care [2]. As healthcare data now extend far beyond EHRs to include imaging, genomics, and signals from wearable devices, these models are increasingly applied to improve patient outcomes [10]. However, most current methods remain limited to a single data type; so they miss important cross-modal information, struggle with incomplete inputs, and often rely on specialized fusion methods that do not scale well to high-dimensional biomedical data. In practice, clinical information is multimodal. Integrating different sources of data is often necessary for accurate diagnosis and effective treatment [1]. Studies show that combining EHRs with imaging or genomics can improve diagnostic accuracy when paired with deep learning methods [11]. At the same time, foundation models—large pretrained systems designed to adapt across many tasks [6] —are becoming increasingly important in healthcare for their ability to generalize across data types and domains [9],[12]. Together, these developments suggest a natural convergence: Multimodal foundation models provide a way to unify heterogeneous clinical data within a single framework, with the potential to improve early disease detection and enable more precise, patient-centered care.
@@ II-B Attention-Based Models
Even so, current multimodal healthcare systems are usually limited. For instance, most models concentrate on single inputs, and leave out other valuable sources like wearables or genomics. Another limitation is that few of them use foundation-style pretraining, where the model is trained broadly across tasks and then reused for specific problems. Without this, adaptability is reduced and each new task often needs a separate model.
This is the gap we target in our work. By combining multimodal integration with foundation-model pretraining, we aim to create patient representations that are broad enough to transfer between tasks yet sensitive to early signals of disease that might otherwise go unnoticed.
@@ III Proposed Framework
We propose a multimodal transformer-based framework to combine diverse biomedical data sources for early illness diagnosis. Unlike unimodal techniques, which handle only one type of input (e.g., EHR or imaging), our architecture combines a variety of clinical data streams, such as electronic health records, medical imaging, genomic sequences, and wearable sensor data. The architecture stresses both flexibility and robustness, allowing predictions even when some modalities are absent, as well as extensibility to new patient data sources that were not available during training. Figure 1 shows a patient-centric multimodal transformer design.
@@ III-A Problem Definition
As not all patients will have complete data across all modalities, our framework is designed to accommodate incomplete records by incorporating modality dropout during training as well as enabling inference on any available subset of $\{X^{ehr},X^{img},X^{gen},X^{sens}\}$, which ensures practical applicability in clinical settings where data coverage is uneven.
@@ III-B Modality-Specific Encoders
Here, raw inputs are structured into feature spaces. For EHR, sequential embeddings capture temporal dependencie, for imaging data are processed through Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs) to extract spatial hierarchies, genomics are encoded with sequence models such as 1D CNNs, while wearable signals are modeled with temporal CNNs or Gated Recurrent Units (GRUs). By design, each of the encoder learns modality-specific representations $h^{m}$ that preserve essential features while normalizing heterogeneous input types.
@@ III-D Training Strategy
During pretraining, the encoders and fusion module learn general multimodal patterns, which helps the model handle noisy data. On the other hand, masked reconstruction encourages it to fill incomplete information, while contrastive learning aligns embeddings from different modalities. In the fine-tuning stage, the network is tailored to the disease detection task by optimizing cross-entropy loss on labeled data. This two-step process strikes a balance between broad general knowledge and task-specific performance.
@@ III-E End-to-End Reasoning Flow
This the chain-of-thought template that depicts the proposed framework’s end-to-end reasoning process to supplement the modality-specific algorithms. This template shows how the whole thing works as a diagnostic workflow. The process starts with raw data preprocessing and modality-specific encoding, then moving on to cross-modal fusion and prediction, uncertainty estimates, explanation creation, and updates driven by feedback.
IV Experimental Design
In order to evaluate the proposed multimodal transformer framework rigorously, we present a comprehensive evaluation strategy that demonstrates clinical relevance, robustness, and feasibility. Although the experiments are conceptual, the design guarantees that the framework can be systematically validated upon its implementation.
IV-A Datasets
The framework is intended to manage a wide range of multimodal patient records. We suggest that an evaluation be conducted on publicly available datasets that encompass complementary types of data across various disease domains:
  • MIMIC-IV: Enables the evaluation of longitudinal modeling and temporal feature extraction by providing structured EHR data, laboratory tests, and clinical notes.
  • UK Biobank: Supports cross-modal integration and generalizability testing across population-level cohorts by incorporating genomic, health record, and imaging data.
  • ADNI: Provides a benchmark for the prediction of neurodegenerative diseases by integrating cognitive assessments, biomarker data, and MRI and PET imaging.
IV-B Baselines and Comparative Approaches
In order to contextualize performance, the proposed framework would be evaluated in comparison to:
  • Unimodal baselines: LSTM/transformer models for EHR sequences, CNN or ViT models for imaging, and gradient-boosted trees for genomics.
  • Multimodal baselines: VAE-based integration, GNN-based patient similarity approaches, and concatenation-based fusion.
IV-C Evaluation Metrics
Evaluation metrics are chosen to indicate the model’s reliability and its clinical significance.
  • AUROC: Evaluates the model’s ability to correctly classify patients based on their specific diseases.
  • AUPRC: Highlights the importance of performance on positive cases to mitigate class imbalance, which is essential for rare conditions.
  • F1-Score and Accuracy: Articulate the quality of classification in a manner that is both clear and understandable.
  • Calibration Error: Assesses the reliability of predicted probabilities, which is essential for clinical decision support.
IV-D Proposed Analyses
To comprehensively validate the framework, we recommend performing a range of analyses:
  • Multimodal vs Unimodal Approaches: To quantify the benefits of integrating multiple data types, each modality will be evaluated separately and in combination.
  • Lack of modality robustness: The model’s capacity to handle incomplete records will be demonstrated by systematically excluding one or more modalities.
  • Ablation Research: The removal of components such as modality-specific encoders, cross-modal attention, or pretraining objectives will explicate their individual contributions.
  • Tests of Generalizability: The potential of transfer learning and its applicability across diseases will be evaluated by utilizing pretrained representations on diverse datasets or patient populations.
V Discussion
Our proposed multimodal transformer framework offers a structured approach to the integration of heterogeneous clinical data in order to facilitate the early detection of diseases. The framework can capture intricate relationships that unimodal models fail to capture by encoding EHRs, imaging, genomics, and wearable signals and fusing them through cross-modal attention. This design facilitates interpretability through attention analysis, robustness to missing data, generalizable patient representations obtained through pretraining, and enhanced predictive performance. The following challenges persist: high computational demands, limited publicly available multimodal datasets, variability and quality of data, and integration into clinical workflows. Future directions include the integration of the framework with real-time clinical decision support systems, the exploration of lightweight transformer variants, the development of continual learning and domain adaptation strategies, and the addition of new data sources. In general, this method provides a scalable, interpretable path to the identification of early diseases that are more precise and personalized.
@@ VI-A Oncology: Multimodal Detection of Cancer
The detection of early cancer remains one of the most significant challenges in clinical medicine, as the stage at which the disease is diagnosed has a strong correlation with the outcome. Traditional methods heavily depend on imaging or histopathology; however, these modalities may not be able to detect subtle precancerous or early neoplastic changes when used in isolation. A multimodal foundation model can integrate radiological images, pathology slides, genomic alterations, and structured EHR data, including family history and laboratory results. For instance, longitudinal EHR data may offer patterns of symptom evolution, while genomic biomarkers associated with tumor predisposition could be contextualized with radiographic features of suspicious lesions. The fusion of these complementary signals has the potential to identify malignancies at a preclinical stage, thereby reducing false negatives and enhancing screening specificity. In practice, this system could serve as an auxiliary diagnostic aid, indicating high-risk patients for additional evaluation prior to the complete emergence of clinical symptoms.
@@ VI-B Cardiovascular Disease: Predicting Heart Failure
Despite the fact that cardiovascular disease continues to be one of the most prevalent causes of illness and mortality on a global scale, it remains a significant obstacle to early identification of individuals who are at risk. The majority of contemporary models are predicated on static data from electronic health records, including echocardiographic findings or laboratory results. However, other data including wearable sensor data (heart rate variability, activity levels), imaging from echocardiography or cardiac MRI, and genetic risk scores from sequencing could be combined to further develop a multimodal approach. For instance, imaging indicators of diminished ventricular function and genetic predispositions to arrhythmias could be examined in conjunction with subtle variations in cardiac rhythm detected by wearables. These streams of information could be combined to facilitate real-time monitoring and ongoing risk assessment which will allow clinicians to take preventive measures earlier and reduce hospitalizations and deleterious cardiac events as a result.
VI-C Neurology: Early Detection of Neurodegenerative Disorders
Neurodegenerative diseases like Alzheimer’s and Parkinson’s show up slowly, with early signs of disease appearing years before the symptoms do. The currently used models that only employ scanning or cognitive tests often miss the first signs that a disease is getting worse. But in a multimodal approach, neuroimaging (structural and functional), genomic variants linked to neurodegeneration, longitudinal EHR data recording subtle behavioral or cognitive complaints, and continuous monitoring from wearable devices that track motor activity could all be a apart of. Combining these different signs, the model will able to find changes that were happening behind the scenes before they became noticeable as a loss of cognitive function. For example, sleep problems caused by wearables may work well with early imaging signs of hippocampal atrophy to make a full risk score. This makes it possible for patients to be quickly enrolled in early-stage interventions or clinical studies, which increases the chances of finding treatments that change the course of the disease.

### v2

@@ Abstract
Healthcare data now span EHRs, medical imaging, genomics, and wearable sensors, but most diagnostic models still process these modalities in isolation. This limits their ability to capture early, cross-modal disease signatures. This paper introduces a multimodal foundation model built on a transformer architecture that integrates heterogeneous clinical data through modality-specific encoders and cross-modal attention. Each modality is mapped into a shared latent space and fused using multi-head attention with residual normalization. We implement the framework using a multimodal dataset that simulates early-stage disease patterns across EHR sequences, imaging patches, genomic profiles, and wearable signals, including missing-modality scenarios and label noise. The model is trained using supervised classification together with self-supervised reconstruction and contrastive alignment to improve robustness. Experimental evaluation demonstrates strong performance in early-detection settings, with stable classification metrics, reliable uncertainty estimates, and interpretable attention patterns. The approach moves toward a flexible, pretrain-and-fine-tune foundation model that supports precision diagnostics, handles incomplete inputs, and improves early disease detection across oncology, cardiology, and neurology applications.
@@ I Introduction
The increasing availability of heterogeneous patient data creates both analytical challenges and significant opportunities for data-driven healthcare. Modern clinical data include diverse modalities such as medical imaging, longitudinal EHR records and wearable sensor streams, auditory and video data, unstructured clinical text, and molecular measurements including genomics and proteomics [1]. EHR systems further provide detailed longitudinal patient histories across encounters, capturing diagnoses, medications, and procedures that support temporal modeling of clinical trajectories [2, 3]. As precision medicine becomes increasingly central to clinical practice [4], effective clinical decision-making depends on synthesizing these heterogeneous information sources. Yet most predictive models rely on a single modality, limiting their ability to capture cross-modal dependencies and reducing their potential for patient-centric insights.
Recent advances in deep learning architectures have facilitated the learning of complex, non-linear relationships across heterogeneous data modalities.For instance, attention-based architectures can dynamically weight the most relevant signals and, when pretrained on large heterogeneous datasets, learn generalizable representations for downstream clinical applications [5]. This foundation-model paradigm, established in NLP and computer vision [6], motivates multimodal approaches to early disease detection. At the same time, transformer architectures have shown very strong performance in large-scale pretraining, learning contextual representations transferable to many downstream tasks [7]. Foundation models extend such benefits by offering broad generalization and prompting capabilities [8]. In light of these trends, we suggest a transformer-based multimodal foundation model that combines data from electronic health records (EHR), imaging, genomics, and wearable sensors. In this framework, cross-modal attention lets the model see how different modalities are related to each other and then lets it learn from a wide range of tasks, resulting in representations that can be used for many different patients. This unified framework seeks to enhance integrated diagnostic tools that facilitate precision medicine by amalgamating diverse clinical data streams within a singular computational architecture.
@@ II-A Multimodal Integration
The wide use of EHR systems has led to the development of predictive models that can support clinical care [2]. As healthcare data now extend far beyond EHRs to include imaging, genomics, and signals from wearable devices, these models are increasingly applied to improve patient outcomes [9]. However, most current methods remain limited to a single data type; so they miss important cross-modal information, struggle with incomplete inputs, and often rely on specialized fusion methods that do not scale well to high-dimensional biomedical data. In practice, clinical information is multimodal. Integrating different sources of data is often necessary for accurate diagnosis and effective treatment [1]. Studies show that combining EHRs with imaging or genomics can improve diagnostic accuracy when paired with deep learning methods [10]. At the same time, foundation models—large pretrained systems designed to adapt across many tasks [6] —are becoming increasingly important in healthcare for their ability to generalize across data types and domains [8],[11]. Together, these developments suggest a natural convergence: Multimodal foundation models provide a way to improve early disease detection and enable more precise, patient-centered care.
@@ II-B Attention-Based Models
We address the gaps of single-modality modeling, missing cross-modal fusion, and absent foundation-style pretraining by building transferable patient representations that capture subtle early disease patterns.
@@ III Proposed Framework
We propose a multimodal transformer-based framework to combine diverse biomedical data sources for early illness diagnosis. Unlike unimodal techniques, which handle only one type of input, our architecture combines a variety of clinical data streams, such as EHRs, medical imaging, genomic sequences, and wearable sensor data. The architecture stresses both flexibility and robustness, allowing predictions even when some modalities are absent, as well as extensibility to new patient data sources that were not available during training. Figure 1 shows a patient-centric multimodal transformer design.
@@ III-A Problem Definition
As not all patients will have complete data across all modalities, our framework is designed to accommodate incomplete records by incorporating modality dropout during training as well as enabling inference on any available subset of $\{X^{ehr},X^{img},X^{gen},X^{sens}\}$¿ It also ensures practical applicability in clinical settings where data coverage is uneven.
@@ III-B Modality-Specific Encoders
Here, raw inputs are grouped together using feature spaces. Sequential embeddings are used to capture temporal dependencies for electronic health records (EHR); Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs) are used to extract spatial hierarchies from imaging data; sequence models such as 1D CNNs are used to encode the data for genomics; and temporal CNNs or Gated Recurrent Units (GRUs) are used to model the data for wearable signals. Each encoder learns modality-specific representations $h^{m}$ that ensure that all input types are identical while preserving significant features.
@@ III-D Training Strategy
The encoders and fusion module learn general multimodal patterns during pretraining. This helps the model deal with data that is noisy. On the other hand, masked reconstruction pushes it to fill in missing information, while contrastive learning makes embeddings from different modalities fit together. During the fine-tuning stage, the network is made to work better at finding diseases by optimizing cross-entropy loss on labeled data. This two-step process finds a middle ground between having a lot of general knowledge and doing well on specific tasks.
@@ III-E End-to-End Reasoning Flow
This the chain-of-thought template that depicts the proposed framework’s end-to-end reasoning process to supplement the modality-specific algorithms. This template shows how the whole thing works as a diagnostic workflow. The process starts with raw data preprocessing and modality-specific encoding, then moving on to cross-modal fusion and prediction, uncertainty estimates, explanation creation, and updates driven by feedback. This approach aligns with prior provenance-aware digital-twin systems that integrate explainability and lifecycle traceability for clinical AI [17].
IV Multimodal Model Design and Learning Framework
Our proposed model combines different types of clinical data into a single latent space and uses both supervised and self-supervised goals to make it more robust when the data is noisy or missing.
IV-A Multimodal Data
We simulate a multimodal dataset comprising four heterogeneous modalities: (1) temporal EHR sequences, (2) $32\times 32$ imaging patches, (3) 500-dimensional genomic profiles, and (4) 3-channel wearable sensor time-series. Each of the modality we used is generated with structured stochasticity to approximate early-stage disease signatures. Positive cases contain weak, partially overlapping signals (e.g., soft EHR trends, mild genomic up-regulation, or subtle imaging lesions), while negative cases occasionally exhibit pseudo-pathological artifacts to induce class overlap.
At the same time, to mimic real-world uncertainty, we introduce $10\%$ label noise and apply a $30\%$ random missing-modality dropout, which produces patient records containing arbitrary subsets of modalities. Table I summarizes key data-generation parameters.
IV-B Model Architecture
Each modality is encoded via a dedicated neural encoder mapping raw inputs to a shared 64-dimensional latent space: a GRU for EHR, a two-stage CNN for imaging, a two-layer MLP for genomics, and a temporal convolutional network for wearable signals. Formally,
$h_{m}=f_{m}(X_{m}),\quad h_{m}\in\mathbb{R}^{64},\quad m\in\{\text{ehr,img,gen,wear}\}.$
The embeddings are stacked and passed through a multihead self-attention transformer that performs cross-modal message passing:
$Z=\text{Transformer}\left([h_{\text{ehr}},h_{\text{img}},h_{\text{gen}},h_{\text{wear}}]\right),$
followed by mean pooling to obtain the fused representation $z\in\mathbb{R}^{64}$. Attention weights are retained for interpretability.
Self-supervised components include: (1) modality masking, where one or more $X_{m}$ are zeroed-out during training; (2) reconstruction decoders $g_{m}(z)$ trained via $\ell_{2}$ losses; and (3) a CLIP-style contrastive alignment objective applied to cross-modal pairs:
$\mathcal{L}_{\text{con}}=\frac{1}{2}\left[\text{CE}(S,I)+\text{CE}(S^{\top},I)\right],\quad S_{ij}=\frac{\tilde{h}_{\text{ehr},i}^{\top}\tilde{h}_{\text{img},j}}{\tau}.$
A two-layer MLP with Monte Carlo dropout produces final predictions and provides uncertainty estimates via stochastic forward passes.
IV-C Training Procedure
Training together lowers the supervised cross-entropy loss, the reconstruction loss over masked modalities, and the contrastive alignment loss:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\alpha\,\mathcal{L}_{\text{recon}}+\beta\,\mathcal{L}_{\text{con}},\quad\alpha=\beta=0.1.$
In this step, we train for five epochs using the Adam optimizer (learning rate $10^{-3}$, batch size 16); all of the components, including data simulation, encoders, fusion transformer, decoders, contrastive module, and classifier, are implemented in PyTorch to ensure reproducibility.
TABLE I: Key parameters for data generation and model configuration.
Component Parameter Value
EHR sequence length $T=10$, $d=12$
Imaging resolution $32\times 32$
Genomics dimension 500500 features
Wearable timeseries $T=100$, 33 channels
Label noise $10\%$ flip probability
Missing modalities $30\%$ dropout
Embedding dimension 6464
Fusion transformer 4-head MHA, 128-unit FFN
Classifier 2-layer MLP + dropout
Optimizer Adam ($10^{-3}$ LR)
Batch size 16
Epochs 5
V Results & Discussion
V-A Results
Table II provides a brief summary of the performance of the suggested multimodal foundation model on the held-out test set. The model was trained throughout five epochs, and the optimization was steady as the total loss decreased from 63.42 in the first epoch to 43.56 in the final epoch. Quantitative assessment indicates that the model demonstrates robust predictive performance in the presence of noise and incomplete modalities, as evidenced by balanced classification metrics and elevated area-based scores.
TABLE II: Performance of the multimodal model on the held-out test set.
Metric Value
Training batches 50
Test batches 13
Final training loss (Epoch 5) 43.5574
Accuracy 0.84
Precision 0.8378
Recall 0.8692
F1-Score 0.8532
AUROC 0.8996
AUPRC 0.9059
V-B Discussion
The ROC curve (Fig. 1) has a smooth, concave shape and an AUROC of 0.900. This means that the fused representation works well to separate cases of early disease from cases of no disease, even when some modalities are missing or noisy. The precision–recall curve (Fig. 3) also showcases high sensitivity, with an AUPRC of 0.906 substantially exceeding the class prior baseline. This suggests that the model maintains reliable precision across a wide range of recall thresholds, which is essential for early detection tasks where false negatives are costly.
Figure 2: ROC curve with 95% confidence interval (AUROC = 0.900).
Figure 3: Precision–recall curve with 95% confidence interval (AUPRC = 0.906).
Calibration analysis (Fig. 4) indicates that predicted probabilities align closely with observed frequencies, with the smoothed trend closely mirroring the ideal diagonal. This means that the uncertainty estimates made by Monte Carlo dropout are very close to the real risk, which means they are good for choosing thresholds and making decisions later on. These trust and transparency issues are similar to blockchain-enabled explainable AI frameworks that make sure clinical AI outputs can be checked and verified [18].
Figure 4: Calibration curve showing alignment between predicted probabilities and observed event frequencies.
When looked at together, the results show that cross-modal fusion, reconstruction-based self-supervision, and contrastive alignment all make systems more robust to clinical signals that are incomplete or unclear. Although the performance is promising, subsequent efforts will authenticate the framework using actual multimodal patient datasets and investigate the scalability to higher-capacity encoders for more intricate clinical modalities.
@@ VI-A Oncology: Multimodal Detection of Cancer
Doctors still have a hard time finding cancer early because the stage of diagnosis has a direct effect on how long a person lives. Imaging or histopathology are usually used in traditional workflows, but these methods might not be able to find small precancerous or early neoplastic changes on their own. A multimodal foundation model can look at structured electronic health record (EHR) data, pathology slides, genomic variants, and radiological images all at once. Things like your family history and lab results are included. Longitudinal EHR patterns may signify symptom advancement, whereas genomic biomarkers of tumor susceptibility can be assessed in conjunction with radiographic features of alarming lesions. When you put these two signals together, it’s easier to find early-stage cancers. There are fewer false negatives, and the screening is more accurate.
@@ VI-B Cardiovascular Disease: Predicting Heart Failure
Heart disease is still one of the top causes of death and illness around the world. But it’s still very hard to find people who are at high risk early on. Current models often rely on static EHR snapshots, such as echocardiographic results or laboratory values. A multimodal approach can use information from wearable sensors (like heart rate variability and activity levels), cardiac imaging (like echocardiography and MRI), and genomic risk scores to learn more about how a person’s body and genes work together. Using these methods together lets you keep checking for risks and encourages quick medical action, which could keep people from going to the hospital and having bad heart outcomes.
VI-C Neurology: Detection of Neurodegenerative Disorders
Current models for cardiovascular disease risk prediction often depend on static electronic health record snapshots, which include echocardiographic results or laboratory values. A multimodal approach can use data from wearable sensors (like heart rate variability and activity levels), cardiac imaging (like echocardiography and MRI), and genomic risk scores to give us a better idea of how a person’s body works and what their genes are made of. People can keep checking their risk and get medical help right away by using these methods together. This could keep them from going to the hospital and having heart problems.